Lauren plans to deposit $8000 into a bank account at the beginning of next month and $250/month into the same account at the end of that month and at the end of each subsequent month for the next 4 years. If her bank pays interest at a rate of 5%/year compounded monthly, how much will Lauren have in her account at the end of 4 years?
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+5/(12*100))^12-1)*100
Effective Annual Rate% = 5.1162
Future value = present value*(1+ rate)^time
Future value = 8000*(1+0.051162)^4
Future value = 9767.17
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 250*(((1+ 5/1200)^(4*12)-1)/(5/1200))
FV = 13253.72

total FV =9767.17+13253.72

=

23020.89